OpenStudy (anonymous):

I have he a question on this one equation that someone sovled. An isosceles triangle in which angles B and C have the same measure. Find the measure of each angle whose degree is represented with variables. A= 4x + 2y B= 9y - 2 C= 10x + 10 We know B = C,so we know 10x + 10 = 9y - 2 thus x = (9y -12)/10 and A + B + C = 180 14x + 11y + 8 = 180 14 (9y - 12)/10 + 11y = 172 126y - 168 + 110y = 1720 236y = 1888 y = 8 x = (72 - 12)/10 = 6 So B = 70 C = 70 A = 40 On Equation x = (9y -12)/10 how did someone get 14 on this equation: 14 (9y - 12)/10 + 11y = 172

OpenStudy (anonymous):

14x + 11y + 8 = 180 Replace x above by (9y -12)/10 from the equation "thus x = (9y -12)/10" three lines back. 14 (9 y - 12)/10 + 11 y = 172 Multiply each side by 10 14 (9 y - 12) + 110 y =1720 Expand the left side. -168 + 236 y = 1720 Move -168 to the RHS 236 y=1720+168 236 y=1888 y=8

OpenStudy (anonymous):

Attached is a two step Mathematica solution.